40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1458     Entry time: Wed Apr 8 02:47:42 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking status 
This is a summary of activities in the last few nights, although there is not much progress.

The attachment 1 and 2 show the CARM and DARM responses around 3.8kHz at different arm power levels.
The CARM error signal was PO_DC and the DARM error signal was AS2Q.
The excitations were both applied to the ETMs (I temporarily modified the output matrix so that the unsed XARM filter bank can be used to excite CARM and DARM).
DARM and CARM show very similar behavior as the power goes up.

The third attachment shows transfer functions to various signals from CARM and DARM excitations (ETMs).
Though the plot contains many curves, look at PO_DC curves (green and black).
PO_DC is used as CARM error signal but it has a larger response to DARM than CARM (by 10dB or so).
This is not good.

Although the 3.8kHz problem still exists, tonight I was able to go up to arm power = 80 a couple of times, where we are ready to hand off from PO_DC to the RF CARM signal. The hand off failed. I'm now optimizing the hand off gain, but it is difficult because the interferometer is unstable at this power level.
Attachment 1: CARM_TFs.pdf  7 kB  | Hide | Hide all | Show all
CARM_TFs.pdf
Attachment 2: DARM_TFs.pdf  6 kB  | Show | Hide all | Show all
Attachment 3: DARM-CARM-Coupling.pdf  8 kB  | Hide | Hide all | Show all
DARM-CARM-Coupling.pdf
ELOG V3.1.3-